Oxford Materials IT Committee Meeting

TRINTY TERM 2001
2 May 2001 2pm 21BR committee room

AGENDA

  • 1 Apologies for absence
  • 3 Sabbatical
  • 3 Matters arising from the minutes
  • 4 University IT Users Group - matters for report (AC)
  • 5 IT/AV Committee
  • 6 IT/AV Staffing Levels
  • 7 IT/AV Documentation
  • 8 Departmental scanning/printing facilities
  • 9 Migration of network connections from thin Ethernet to UTP
  • 10 Update on Status of Department Ethernet
  • 11 Report upon a computer attack.
  • 12 Any Other Business
  • 13 Next Meeting


MINUTES

Minutes of the IT/AV Committee held at 2.00pm on 17 October 2001

Present: GDB, Alex Cock, MB, AC, PJW

1. Apologies: AMcK

2. Sabatical: JTC

3. Matters arising from the minutes

  • 2(b) MB still to circulate notes about Sophos Automatic updating for Windows 95, 98, ME. Action: MB
  • 2(b) AC reiterated that every single piece of site-license software needs an order number, even if it is free.
  • 2(c) Development of Intranet in conjunction with General Office is on hold. (PJW/Michelle)
  • 2(e) Department policy on printing hardcopies of the Research in Progress document is to be revisited, but this should wait until the new administrator arrives. (AMcK)
  • 7 The migration of POP users to Herald appears to have gone smoothly. The committee would like to thank Matt and Alex for all their help to Department members during the transition.
  • 8 Original plans for issuing portable computers to all graduate students have now been abandoned in the light of the high cost involved and lack of any sponsors being identified.

4. Oxford University IT Users Group Meeting http://www.ict.ox.ac.uk/oxford/groups/ouitug/

a) Sable will be taken out of service on 31st July 2001. AC will circulate advice to members of department. Those who only use sable for email should migrate to herald, whilst those who need access to UNIX should register for and account on ermine, unless they are using MML facilities. Action: AC

b) There will be an IT Support Staff conference on 1st June. Alex and MB and PJW to attend and need to register before 31st May. Action: Alex Cock / MB / PJW .

c) The university network support service offers Sophos auto-updating at £10/box/year. However we have now made arrangements to provide this service in-house (except problem with Windows ME) so there is no need to out-source. Guidance notes are soon to be distributed to Windows users. Action: MB / Alex Cock

5. IT/AV Committee

a) AC reiterated what he felt to be the role of the committee within the structure of the Department, and this was discussed. It was considered to be useful if AC produced a formal Terms of Reference document.

b) It was noted that WRCB and SDK have left the department. The committee wishes to thank them for all their help during their time in the department. Is was agreed that apart from the new administrator, who might wish to join the committee, a member of the EM group should be asked to join since they are significant users of IT facilities. Action: AC

6. IT/AV Staffing Levels

Further to the Review of IT/AV and the request for three new posts (Senior IT Officer, IT Technician, AV Technician) AC reported that the RIC decided that due to a shortage of funds not all posts could be funded. Due to the departure of SDK, support for the MML was top priority and hence the Senior IT Manager post (RSII) is to be funded. (Currently PJW is providing support for the MML as an interim measure.) RIC committed to increase IT funding though grant applications and to reconsider the other posts when the funding situation of the Department becomes clearer over the summer vacation. AC noted his disappointment in this outcome.

A draft job description for the Senior IT Officer was tabled and discussed. AC is to discuss the job description further with Peter Dobbs (Administrator, Physics) before submitting to RIC. Action: AC

7. IT/AV Documentation

Further to the discussion of the terms of reference of the committee, a series of documents were suggested to form the basis of this departments IT/AV committee.

  • Committee Terms of Reference Action: AC
  • Department IT Rules Action: AC / GDB
  • Computer Recycling/Reuse policy Action: AC / GDB
  • Potential for DHCP service provision Action: Alex Cock
  • Instructions for automatic anti-virus protection Action: MB
  • Network printing instruction Action: MB
  • Network status Action: MB
  • Departmental IT/AV facilities Action: All / PJW
  • MML guidance notes Action: PJW

8. Departmental scanning/printing facilities

AC and AMcK have assessed the hardware requirements for the new Digital Darkroom facility (located beside photographic). On the basis of the need for high speed, high quality, colour transparencies, a solid wax printer has been selected. There will also be an A3 colour laser printer for lower cost paper printing and poster production. Total cost for two printers and two computers (1PC 1 Mac) will be about £12k inc vat. Access to the room will be via University Card swipe-access same as the EM area. There will be some additional cost (£1k max) for installation of the card entry system.

High speed general purpose printing is also becoming available through networking of the departmental photocopiers. Several actions are required to make these facilities fully functional.

  • General office photocopier is set-up already and has been used by PJW.
  • Top floor HR photocopier to be networked Action:GDB
  • Charge-system to be implemented and checked Action:GDB/PJW
  • Printer driver set-up instructions to be prepared for Mac and PC Action: MB
  • System to be checked thoroughly before making generally available Action: All

9. Migration of network connections from thin Ethernet to UTP

Alex and MB now use Transcend software to locate computers on the network. A list of people still using thin Ethernet in the HR and HB has been compiled and people will be contacted to help them move to UTP connections. The aim is to move all HR staff to UTP network as soon as possible since this service is at risk. HB will be the next area to move entirely to UTP.

10. Update on Status of Department Ethernet

Programme of conversion to UTP continues.

HR - done (except S16? and library?)

HB - done (except EM area and Connie)

ETB - photographic and electronics recently complete. Line to be added in LR7 but other areas not currently a problem so are not a priority

PR - First and second floors - 12 done and 13 to be done during current refurbishment. Ground floor and basement still to do. No schedule.

21BR - MML and Polymer/Biomaterials areas done. Top offices on back staircase fixed temporarily prior to planned refurbishment. Front offices done. Need to transfer users and strip out thin ethernet. BC labs and offices above are experiencing problems. Need to check/shorten segment. Hopefully this can be improved with other changes in the building. Action: MB

Overall, 21BR requires the most attention at the moment, and generally needs mapping so this is a priority area for expenditure. Potential to rewire when new distribution boards and trunking is installed later this year.

GDB to prepare updated cost estimate for completion of programme. Action: GDB

12. Report upon a computer attack.

Just before Easter one computer is this Department was subject to a flood attack which amounted to 35 times to traffic of the rest of the university combined. The particular computer seems to have been targeted because it was running an IRC (chat room) server. AC will circulate a general note on the risks of this type of attack pointing out its effect upon other Department and University users, and stressing the importance of computer security. Action: AC

13. Any Other Business

a) PJW to formally become a member of IT support staff and join itss-annouce and itss-discuss mailing lists. Action: PJW

b) Computer recycling procedure to be clarified - specifically for the old MML hardware. Action: PJW

14. Next Meeting:

Wednesday 17 October 2001, 2.15pm. However, it was decided that an interim meeting might be held during the summer, once the Senior IT Officer had been appointed.

Meeting Closed 5.30pm